OpenStudy (anonymous):

Let θ be an angle in quadrant II such that cos θ = -12/13. Find the exact value of csc θ and tan θ .

OpenStudy (anonymous):

\[\sin \theta =\sqrt{1-\cos^2\theta}\]

OpenStudy (anonymous):

5/13

OpenStudy (anonymous):

\[\csc \theta =13/5\]

OpenStudy (amistre64):

5-12-13 ;)

OpenStudy (anonymous):

\[\tan \theta=-5/12\]